# EarthquakeNPP: Benchmark Datasets for Earthquake Forecasting with Neural Point Processes

### v1

@@ 5.1 Number Test
5.2 Spatial Test
@@ 5.3 Magnitude Test
5.4 Evaluating Multiple Forecasting Periods
@@
Savran et al. (2020) describe how to assess a model’s performance across the multiple days in the testing period. By construction, quantile scores over multiple periods should be uniformly distributed if the model is the data generator (Gneiting & Katzfuss, 2014). Therefore comparing quantile scores against standard uniform quantiles (y = x), highlights discrepancies between the observed data and the forecast. Additional statements can be made about over-prediction or under-prediction of each test statistic (quantile curves above/bellow y=x respectively). The Kolmogorov-Smirnov (KS) statistic then quantifies the degree of difference to the uniform distribution for each of the tests.
Further documentation of how to perform the CSEP evaluation procedure can be found on the platform, where we demonstrate the procedure for the ETAS model. Table 1 reports the benchmark performance scores taken from the quantile plots in Appendix C. The performance of ETAS is higher for the more typical higher magnitude catalogs such as ComCat and SCEDC, whereas it performs worse at the lower magnitude catalogs of QTM_SanJac, QTM_SaltonSea and White. Spatial prediction is consistently the best performing component of the ETAS forecast, whereas earthquake numbers are overpredicted by the model and earthquake magnitudes are generally not well predicted (Figure 9) . All results indicate significant room for improvement beyond the predictive performance of the ETAS model.
Table 1: CSEP consistency tests evaluate the calibration of all daily ETAS forecasts on EarthquakeNPP datasets. A test is performed at the $\alpha=0.05$ significance level on each day in the testing period. The pass rate indicates the success of ETAS across all testing days. By construction quantile scores of the tests should be uniformly distributed if the model is the data generator. The KS-Statistic reports the difference of the quantile distribution to uniform, taken from the quantile plots in Appendix C.
Dataset Number Test Spatial Test Magnitude Test
Pass Rate KS-Statistic Pass Rate KS-Statistic Pass Rate KS-Statistic
ComCat 62.3% 0.392 85.3% 0.128 75.3% 0.318
SCEDC 74.4% 0.161 87.5% 0.123 80.5% 0.153
QTM_SanJac 59.2% 0.461 96.7% 0.145 66.2% 0.406
QTM_SaltonSea 54.2% 0.441 82.1% 0.216 79.0% 0.311
White 17.1% 0.750 98.0% 0.373 25.0% 0.741
6 Discussion and Conclusion
We introduce the EarthquakeNPP datasets to facilitate the benchmarking of NPPs against a community-endorsed ETAS model for earthquake forecasting. These datasets cover various regions of California, representing typical forecasting zones and the data commonly available to forecast issuers. Several datasets use modern methods of detection, which enables the inclusion of much smaller magnitude earthquakes.
In a benchmarking experiment, we compared three NPP models against ETAS and a baseline Poisson process. None of the NPP models outperformed ETAS, indicating that current NPP implementations are not yet suitable for operational earthquake forecasting. ETAS explicitly defines how larger earthquake magnitudes increase the likelihood of future earthquakes in both time and space, with an empirical relationship derived from observational studies (Utsu & Seki, 1955; Utsu, 1970). Since the NPPs lack any direct dependence on magnitudes, this is the likely cause for their performance relative to ETAS. Future implementations should exploit this additional feature for improved temporal and spatial performance. Encouragingly, the comparable temporal performance to ETAS without this additional feature suggests that incorporating magnitude dependence would enhance NPP performance beyond that of ETAS.
EarthquakeNPP supports the earthquake forecast evaluation procedure developed by the Collaboratory for the Study of Earthquake Predictability (CSEP). The procedure replicates how earthquakes forecasts are generated in an operational setting, requiring models to simulate many repeat event sequences over a day-long forecast horizon. Benchmark performance for the ETAS model enables future comparison of NPPs that are implemented for this procedure and enables their promotion to the fully prospective CSEP experiments. Notably, this procedure allows the evaluation of generative NPP models without explicit likelihoods (Yuan et al., 2023; Li et al., ), by assessing their performance over the full trajectory of future events. Probabilistic seismic hazard analysis (PSHA) requires long-term prediction beyond the next-event (Ebrahimian et al., 2014; Gerstenberger et al., 2014), therefore this approach also offers stakeholders a more comprehensive understanding of earthquake hazard than metrics focused on predicting the next event (e.g. RMSE). The procedure also follows the recommendation by Shchur et al. (2021) to move away from next-event point prediction for NPPs.
The EarthquakeNPP datasets, available at https://anonymous.4open.science/r/EarthquakeNPP-2D51, provide a platform for future NPP developments to be benchmarked against these initial results. The platform is under ongoing development and in the future will see the direct comparison of emerging and other existing models models developed within the seismology community, as well as an expansion of datasets included to other seismically active global regions. Successful NPP models on these datasets, for both log-likelihood and CSEP metrics, will be directly impactful to stakeholders in seismology, ultimately enabling their integration into operational earthquake forecasting by government agencies.
Appendix A Earthquake Catalog Data
A.1 Earthquake Catalog Generation
Seismometers and Seismic Networks. A seismometer is an instrument that detects and records the vibrations caused by seismic waves (Stein & Wysession, 2009; Shearer, 2019). It consists of a sensor to detect ground motion and a recording system to log three-dimensional ground motion over time, typically vertical and horizontal velocities. Seismic networks, comprising multiple seismometers, monitor seismic activity at regional, national or global scales (see, e.g., (Woessner et al., 2010) and references therein). High-density networks with modern, sensitive equipment provide more detailed and accurate data, enhancing the ability to detect and analyse smaller and more distant earthquakes.
From Waveforms to Phase Picking. The process of converting raw continuous seismic waveforms into useful earthquake data begins with phase picking, which identifies the arrival times of the primary (P) and secondary (S) waves of an earthquake. Historically, this was done manually, but now automated algorithms, such as the STA/LTA algorithm, detect wave arrivals by analyzing signal amplitude changes (Allen, 1982). Recent algorithms, such as machine learning classifiers (e.g. Zhu & Beroza, 2019; Lapins et al., 2021) and template-matching (e.g. Ross et al., 2019), can process much higher volumes of data efficiently and are often able to detect events of much smaller magnitudes.
Earthquake Association and Location After phase picking, the next step is to associate phases from different seismometers with the same earthquake. Simple algorithms require at least four phase arrivals to be detected on different stations within a short time interval to declare an event. Once phases are associated, location estimation determines the earthquake’s hypocenter and origin time by minimizing travel-time residuals using linearized or global inversion algorithms (Thurber, 1985; Lomax et al., 2000). Given the potential for misidentified or mis-associated phase arrivals due to low signal-to-noise of small events or the near-simultaneous occurrence during very active aftershock sequences, an automated system typically first picks arrival times and determines a preliminary location, which is subsequently reviewed by a seismologist (e.g. Woessner et al., 2010, and references therein). Locations are typically reported as the geographical coordinates and depths where earthquakes first nucleated (hypocenters), although some catalogs report the centroid location, a central measure of the extended earthquake rupture.
Earthquake Magnitude Calculation The magnitude of an earthquake quantifies the energy released at the source and was originally defined in the seminal paper by Richter (1935). The original definition, now referred to as the local magnitude (ML), is calculated from the logarithm of the amplitude of waves recorded by seismometers. This scale, however, "saturates" at higher magnitudes, meaning it underestimates magnitudes for various reasons. This led to introduction of the moment magnitude scale (Mw) (Hanks & Kanamori, 1979), which computes the magnitude based on the estimated seismic moment $M_{0}$, which can be related to the physical rupture process via
$M_{0}=\text{rigidity}\times\text{rupture area}\times\text{slip},$ (16)
where rigidity is a mechanical property of the rock along the fault, rupture area is the area of the fault that slipped, and slip is the distance the fault moved. Mw is determined seismologically via a spectral fitting process to the earthquake waveforms. In practice, it can be challenging to use a single magnitude scale for a broad range of magnitudes, therefore a range of scales may be present within a single catalog, and approximate magnitude conversion equations may be used to homogenize the scales (e.g. Herrmann & Marzocchi, 2021, and references therein).
A.2 Earthquake Catalog Completeness
All of the EarthquakeNPP datasets are made publicly available by their respective data centers in raw format. However, constructing a suitable retrospective forecasting experiment from this raw data requires appropriate pre-processing. This typically involves truncating the dataset above a magnitude threshold $M_{\text{cut}}$ and within a target spatial region to address incomplete data, known as catalog completeness $M_{c}$ (e.g., Mignan et al., 2011; Mignan & Woessner, 2012).
There are several reasons why an earthquake may not be detected by a seismic network. Small events may be indistinguishable from noise at a single station, or insufficiently corroborated across multiple stations. Another significant cause of missing events occurs during the aftershock sequence of large earthquakes, when the seismicty rate is high (Kagan & Knopoff, 1987; Hainzl, 2022). Human or algorithmic detection abilities are hampered when numerous events occur in quick succession, e.g. when phase arrivals of different events overlap at different stations or the amplitudes of small events are swamped by those of large events. Since catalog incompleteness increases for lower magnitude events, typically the task is to find the value $M_{c}$ above which there is approximately $100\%$ detection probability. Choosing a truncation threshold $M_{\text{cut}}$ that is too high removes usable data. Where NPPs have demonstrated an ability to perform well with incomplete data (Stockman et al., 2023), typically a threshold below the completeness biases classical models such as ETAS (Seif et al., 2017). Seismologists often investigate the biases of different magnitude thresholds by performing repeat forecasting experiments for different thresholds (e.g. Mancini et al., 2022; Stockman et al., 2023), which we also facilitate in our datasets.
Typically $M_{c}$ is determined by comparing the raw earthquake catalog to the Gutenberg-Richter law (Gutenberg & Richter, 1936), which states that the distribution of earthquake magnitudes follows an exponential probability density function
$f_{GR}(m)=\beta e^{\beta(m-M_{c})}\ \ :m\geq M_{c}.$ (17)
where $\beta$ is a rate parameter related to the b-value by $\beta=b\log 10$. Histogram-based approaches, such as the simple Maximum Curvature method (Wiemer & Wyss, 2000) as well as many others (e.g. Herrmann & Marzocchi, 2021, and references therein), identify the magnitude at which the observed catalog deviates from this law, indicating incompleteness (See Figure 6b).
In practice, catalog completeness varies in both time and space $M_{c}(t,\mathbf{x})$ (e.g. Schorlemmer & Woessner, 2008). During aftershock sequences, $M_{c}(t)$ can be very high (e.g., Agnew, 2015; Hainzl, 2016b) (See Figure 6a). Thresholding at the maximum value might remove too much data. Instead, modelers either omit particularly incomplete periods during training and testing (Kagan, 1991; Hainzl et al., 2008), model the incompleteness itself (Helmstetter et al., 2006; Werner et al., 2011; Omi et al., 2014; Hainzl, 2016a; b; Mizrahi et al., 2021; Hainzl, 2022), or accept known biases from disregarding this issue (Sornette & Werner, 2005). Spatially, catalogs are less complete farther from the seismic network (Mignan et al., 2011), so the spatial region can be constrained to remove outer, more incomplete areas (See Figure 6c).
Appendix B Additional Benchmark Results
Appendix C CSEP Consistency Tests
Figure 9: Quantile-quantile plots showing the calibration of all daily ETAS forecasts on a) ComCat,b) SCEDC, c) QTM_San_Jac, d) QTM_Salton_Sea, e) White. By construction quantile scores over multiple periods should be uniformly distributed if the model is the data generator. Comparing quantile scores against standard uniform quantiles ($y=x$), highlights discrepancies between the observed data and the forecast. Pass rates of each test are indicated in the legend. The Kolmogorov-Smirnov statistic, quantifies the degree of difference to the uniform distribution.

### v2

@@ E.1 Number (Temporal) Test
E.2 Pseudo-Likelihood Test
The pseudo-likelihood test evaluates the compatibility of a forecast with an observed catalog using an approximation to the space-time point process likelihood.
The test statistic is based on the pseudo-log-likelihood:
$\hat{L}_{\text{obs}}=\sum_{i=1}^{N_{\text{obs}}}\log\hat{\lambda}_{s}(k_{i})-\bar{N},$ (12)
where $\hat{\lambda}_{s}(k_{i})$ is the approximate rate density in the spatial cell of the $i^{\text{th}}$ event, and $\bar{N}$ is the expected number of events.
Each forecast simulation $j$ provides a test statistic
$\hat{L}_{j}=\sum_{i=1}^{N_{j}}\log\hat{\lambda}_{s}(k_{ij})-\bar{N},$ (13)
which is used to build the empirical cumulative distribution $F_{L}$. The quantile score is then computed as
$\gamma_{L}=\mathbb{P}(\hat{L}_{j}\leq\hat{L}_{\text{obs}})=F_{L}(\hat{L}_{\text{obs}}).$ (14)
E.3 Spatial Test
@@ E.4 Magnitude Test
(a) Number test
(b) Spatial test
(c) Magnitude test
@@
E.5 Evaluating Multiple Forecasting Periods
@@
Savran et al. [63] describe how to assess a model’s performance across the multiple days in the testing period (Figure 12). By construction, quantile scores over multiple periods should be uniformly distributed if the model is the data generator [15]. Therefore comparing quantile scores against standard uniform quantiles (y = x), highlights discrepancies between the observed data and the forecast. Additional statements can be made about over-prediction or under-prediction of each test statistic (quantile curves above/bellow y=x respectively). The Kolmogorov-Smirnov (KS) statistic then quantifies the degree of difference to the uniform distribution for each of the tests.
Figure 12: Daily number forecasts from SMASH on the ComCat dataset. (a) Forecasted daily distributions for the number of earthquakes, with green lines indicating days where the observed count falls within the 95% forecast interval, and red lines where the forecast fails. Observed values are marked with dot sizes proportional to the number of earthquakes. (b) Quantile scores from the number test for each day, with red markers indicating failed forecasts. Marker size reflects the number of earthquakes observed on that day. (c) Temporal evolution of observed earthquakes during the testing period, with event magnitudes represented by marker size. (d) Histogram of quantile scores from the number test. Under ideal calibration, scores should follow a uniform distribution. Red bars indicate failed forecasts, and the Kolmogorov–Smirnov (KS) statistic quantifies deviation from uniformity.
Appendix F Further Dataset Figures
F.1 ComCat
Figure 13: Times and magnitudes of events in the ComCat dataset (with key events labeled). The size of the points are plotted on a log scale corresponding to Mw. Auxiliary, training, validation and testing periods are indicated by colour and a further cumulative count of events is indicated in red.
Figure 14: Locations of events in the ComCat dataset, labeled by their partition into auxiliary, training, validation and testing periods.
F.2 SCEDC
Figure 15: Times and magnitudes of events in the SCEDC dataset (with key events labeled). The size of the points are plotted on a log scale corresponding to Mw. Auxiliary, training, validation and testing periods are indicated by colour and a further cumulative count of events is indicated in red.
Figure 16: Locations of events in the SCEDC dataset, labeled by their partition into auxiliary, training, validation and testing periods.
F.3 White
Figure 17: Times and magnitudes of events in the White dataset (with key events labeled). The size of the points are plotted on a log scale corresponding to Mw. Auxiliary, training, validation and testing periods are indicated by colour and a further cumulative count of events is indicated in red.
Figure 18: Locations of events in the White dataset, labeled by their partition into auxiliary, training, validation and testing periods.
F.4 QTM_SanJac
Figure 19: Times and magnitudes of events in the QTM_SanJac dataset. The size of the points are plotted on a log scale corresponding to Mw. Auxiliary, training, validation and testing periods are indicated by colour and a further cumulative count of events is indicated in red.
F.5 QTM_SaltonSea
Figure 20: Times and magnitudes of events in the QTM_SaltonSea dataset. The size of the points are plotted on a log scale corresponding to Mw. Auxiliary, training, validation and testing periods are indicated by colour and a further cumulative count of events is indicated in red.
Figure 21: Locations of events in the QTM_SanJac and QTM_SaltonSea datasets, labeled by their partition into auxiliary, training, validation and testing periods.
Appendix G Error Distributions & Next-event metrics
Figure 22: The distribution of errors ($Y_{\text{obs}}-Y_{\text{pred}}$) for the Normal(0,1)(0,1), Exponential(1)(1), and Pareto(2)(2) distributions. Maximum likelihood estimation is used to fit Normal and Laplace distributions to each error histogram. Normal errors (Normal $\times$ Normal) are best approximated by the Root Mean Square Error (RMSE), while Laplacian errors (Exponential $\times$ Exponential) are best approximated by the Mean Absolute Error (MAE). However, neither RMSE nor MAE effectively capture the errors for the heavy-tailed Pareto distribution.